Nitrogen gas reacts with hydrogen gas to produce ammonia.
N2(g)+3H2(g)→2NH3(g)
How many grams of NH3 can be produced from 10.4 g of H2?
Express your answer with the appropriate units.
1
Expert's answer
2015-10-20T09:06:13-0400
Solution:
ν(H2)=10.4/2=5.2 mol
3 mol H2→2 mol NH3, so ν(NH3)=(5.2/3)*2=3.47 mol
m(NH3)=17*3.47-=58.99 g
Answer: 58.99 grams of NH3.